HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

STATE-OF-THE-ART ELECTRONICS FOR LASER-BASED DETECTION OF^{14}CO_{2}

\underline{A. DETTI}, L. MISCHI, F. CARCIONE, T. CONTI, ppqSense Srl, Viale Ariosto 492/B, 50019 Sesto Fiorentino FI, Italy; A. MONTORI, S. BARTALINI, ppqSense Srl, Viale Ariosto 492/B, 50019 Sesto Fiorentino FI, Italy; CNR - Istituto Nazionale

di Ottica - Largo Fermi 6, 50125 Firenze

Combining advanced electronics for controlling semiconductor Quantum Cascade Lasers (QCLs)^{1} with the most recent spectroscopy techniques allows to get ideal metrological probes to measure extremely low concentrations of rare molecules. The best results have been obtained in the last few years with saturated-absorption cavity ring-down (SCAR)^{2}, a nonlinear cavity ring-down technique, to detect the very elusive ^{14}CO_{2}. Indeed, this radioactive molecule amounts only to about 1 part in 10^{12}, as compared to the overall CO_{2} in the biosphere. Therefore, the sensitivity achieved by SCAR, of a few parts in 10^{15}, provides a wide dynamic range for radiocarbon dioxide detection, challenging the best results provided by Accelerator Mass Spectrometry (AMS)^{3}.

110.1088/0957-0233/25/1/012001 , S. Bartalini and M. S. Vitiello and P. De Natale, Quantum cascade lasers: a versatile source for precise measurements in the mid/far-infrared range, Measurement Science and Technology, 25, 1, 012001, (2013).

210.1103/PhysRevLett.104.11080, Giusfredi, G. and Bartalini, S. and Borri, S. and Cancio, P. and Galli, I. and Mazzotti, D. and De Natale, P., Saturated-Absorption Cavity Ring-Down Spectroscopy, Phys. Rev. Lett. 104, 11, 110801, (2010).

310.1364/OPTICA.3.00038, I. Galli, S. Bartalini, R. Ballerini, M. Barucci, P. Cancio, M. De Pas, G.

Giusfredi, D. Mazzotti, N. Akikusa, and P. De Natale, Spectroscopic detection of radiocarbon dioxide at parts-per-quadrillion sensitivity, Optica 3, 385-388 (2016).

p-number: p257  Submitted on Mon, 03 Jul 2023 16:52:37 +0200